OpenStudy (andriod09):

Quadratic Equation question How do i get from: \[x=-11\pm \frac{121-60}{ 6 }\] to the next part?

OpenStudy (andriod09):

@jwheele1

OpenStudy (jwheele1):

ok

OpenStudy (anonymous):

im working on those problems too im not that good but i would simplify 121-6/ then.. which is?

OpenStudy (jwheele1):

well first of all.....write out two different equations to eliminate that +-

OpenStudy (andriod09):

@cos00155079 you doing the Life of Fred series?

OpenStudy (andriod09):

you don't have to do that @jwheele1

OpenStudy (jwheele1):

i know you dont have to but its easier...its cool we wont do that then

OpenStudy (anonymous):

the what? nooo...

OpenStudy (jwheele1):

121 - 60 = ?

OpenStudy (andriod09):

61

OpenStudy (andriod09):

10

OpenStudy (andriod09):

\[1\] \[-21\]

OpenStudy (anonymous):

why 60/6 if it was a 61?

OpenStudy (andriod09):

ik

OpenStudy (jwheele1):

alright lets see now...

OpenStudy (andriod09):

the equation is: \[x=-11\pm \frac{ \sqrt{121-60} }{ 6 } \]

OpenStudy (jwheele1):

oh lawd....lol

OpenStudy (jwheele1):

sq rt of 121 = 11

OpenStudy (jwheele1):

whats with teh line over the 0?

OpenStudy (andriod09):

its just how it is. not my doing.

OpenStudy (jwheele1):

well...it means that its a repeating number but its usually only after a decimal point...weird.

OpenStudy (jwheele1):

I have no idea what to do with the 60 and the line over it.

OpenStudy (andriod09):

the original equation is: \[3x^{2}+11x+5=0\]

OpenStudy (jwheele1):

any specific way she is wanting you to go about that?

OpenStudy (andriod09):

its a messup about square roots.

OpenStudy (jwheele1):

or can we solve for x any old way?

OpenStudy (andriod09):

yea, using \[ax^{2}+bx+c=0\] and the \[x=-b\pm\frac{ \sqrt{b^{2}-4ac} }{ 2a }\]

OpenStudy (jwheele1):

the QuAdRaTiC ForMulA.....lol

OpenStudy (jwheele1):

ok so 3x^2 + 11x +5 = 0 ax^2 + bx +c

OpenStudy (jwheele1):

a = 3 b = 11 c = 5

OpenStudy (jwheele1):

\[-11\pm \frac{ \sqrt{11^{2}-4(3*5c)} }{ 2(3) }\]

OpenStudy (jwheele1):

does that look right?

OpenStudy (jwheele1):

ignore the c next to the 5, lol